9) Several cars parked in a large parking lot sustain hail damage during a storm. The repair costs for the cars are uniformly distributed between $100 and $600.

A car is selected at random from the lot. Find the probability that the repair costs for the car are less than $405.

Provide an exact answer.
Answer #1

Solution :

Given that,

a = $100

b = $600

P(x < c) = (c - a) / (b - a)

P(x < $405) = (405 - 100) / (600 - 100) =305/500=0.61

probability=0.61